find the slope between (1,7) and (3,3)

Answer by persian52(161)

(1,7),(3,3)
Slope is equal to the change in y over the change in x, or 'rise over run'.
m=(change in y)/(change in x)
The change in x is equal to the difference in x-coordinates (also called run), and the change in y is equal to the difference in y-coordinates (also called rise).
m=(y2-y1)/(x2-x1)
Substitute in the values of x and y into the equation to find the slope.
m=(3-(7))/(3-(1))
Multiply -1 by the 1 inside the parentheses.
m=(3-(7))/(3-1)
Subtract 1 from 3 to get 2.
m=(3-(7))/(2)
Multiply -1 by the 7 inside the parentheses.
m=(3-7)/(2)
Subtract 7 from 3 to get -4.
m=(-4)/(2)
Reduce the expression (4)/(2) by removing a factor of 2 from the numerator and denominator.
m=-(2)
Multiply -1 by the 2 inside the parentheses.
Answer: m=-2